# Degeneration of genus three Ceresa cycles and limit of archimedean height

### 02 Apr 2026 v1

@@ 2.2.1. The case of central irreducible nodal curve
We consider the first degeneration described in subsection 1.3.3: $\mathcal{C}\to\Delta$ is a 1-dimensional family of genus $g$ curves, whose fibre over $0\in\Delta$ is an irreducible nodal curve $C_{0}$, the node $x_{0}$ is obtained by identifying two points $p,q$ in the normalization $\widetilde{C}_{0}$. Furthermore, the cycle pqp-q is torsion. The natural variation of Hodge structure, whose fibre over $t\in\Delta$ is provided by $H^{1}(C_{t},\mathbb{Q})$, has a limit MHS, as we now recall. Let $H_{1}(C_{t},\mathbb{Z})=\langle\alpha_{1},\ldots,\alpha_{g},\beta_{1},\ldots,\beta_{g}\rangle$ be a symplectic basis (similarly $\alpha_{1}^{*},\ldots,\beta_{g}^{*}$ the dual basis in cohomology), and $\alpha_{g}$ the vanishing cycle of the degeneration. Then, since $N^{2}=0$, the monodromy weight filtration has 3 weights and
@@
where the equality in $M_{1}$ follows from Theorem 2.1, since up to a base change, our family admits a semistable degeneration. We want the mixed Hodge structure on $H^{1}_{lim}(C_{t};\mathds{Q})$ to be split, namely we want $H^{1}_{lim}=\oplus\operatorname{Gr}^{M}_{k}H^{1}_{lim}(C_{t};\mathds{Q})$, where
@@
as a split mixed Hodge structure. Thus, to compute $H^{k}_{lim}(JC_{t}:\mathbb{Q})$, it is enough to take the exterior power $\wedge^{k}H^{1}_{lim}(JC_{t};\mathbb{Q})\cong\wedge^{k}H^{1}_{lim}(C_{t};\mathbb{Q})$. It follows that the limit MHS $H^{k}_{lim}(JC_{t};\mathbb{Q})$ has 3 weights with respect to $M$, and the graded pieces are as follows:
$\operatorname{Gr}^{M}_{k-1}H^{k}_{lim}(JC_{t};\mathbb{Q})=H^{k-1}(J\widetilde{C}_{0},\mathds{Q});$
$\operatorname{Gr}^{M}_{k}H^{k}_{lim}(JC_{t};\mathbb{Q})=H^{k-2}(J\widetilde{C}_{0};\mathds{Q}(-1))\oplus H^{k}(J\widetilde{C}_{0};\mathds{Q}))\cong H^{k}(\widetilde{X}_{0};\mathds{Q});$
$\operatorname{Gr}^{M}_{k+1}H^{k}_{lim}(JC_{t};\mathbb{Q})=H^{k-1}(J\widetilde{C}_{0};\mathds{Q}(-1)),$
where the middle isomorphism follows from the projective bundle formula, noting that $\widetilde{X}_{0}$ is $\mathbb{P}^{1}$-bundle over $J\widetilde{C}_{0}$.
@@ 2.2.2. The case of central reducible curve
We are in the second degeneration in 1.3.3: we are considering a 1-dimensional family of curves $\mathcal{C}\to\Delta$, whose fibre over $0\in\Delta$ is a reducible curve $C_{0}=C_{1}\cup C_{2}$ obtained as the union of two irreducible smooth components $C_{j}$ of genus $1$ and g1g-1 respectively, with a separating node ${n}=C_{1}\cap C_{2}$. As before, we first consider the VHS provided by $H^{1}(C_{t};\mathds{Q})$, for $t\in\Delta^{*}$. In this case, $N=0$ and the limit MHS trivially is as follows:
@@ 3.1. Geometric biextension
In case of a biextension of geometric origin as above, the corresponding height $\operatorname{ht}(B_{Z,W})$ equals the archimedean height pairing $\langle Z,W\rangle$ between the two cycles (this is a consequence of [Hai90, Proposition 3.2.13]. See also [Pe06, §5] for further insights.). In fact, using the duality of remark 3.5 one can interpret $\operatorname{ht}(B_{Z,W})$ as $\int_{Z}\mathfrak{g}_{W}$, where ${\mathfrak{g}}_{W}$ is a Green form of logarithmic type for the cycle $W$ that can be shown to exist. For example, in $\mathds{P}^{1}$ any choice of four distinct points $\{p,q,r,s\}$ gives rise to homologically (actually rationally) trivial cycles Z=pqZ=p-q and W=rsW=r-s. In this case, one choice for ${\mathfrak{g}}_{W}$ is the function $\log\Big|\frac{t-r}{t-s}\Big|^{2}$ that has logarithmic singularities on $|W|=\{r,s\}$. The height is given by the logarithm of the cross-ratio of the four points.
@@ 3.2. Abstract degeneration of biextension variations
Unfortunately, this function does not always extend continuously to $\Delta$. There is a height jump $\mu\in\mathds{Q}$, such that the amended function
@@
with respect to the original weight filtration $W$. Hence the components $N_{-2}$ are $\delta_{M,-2}$ are easy to extract: $N_{-2}\colon\mathds{Q}(0)\to\mathds{Q}(1)$ and $\delta_{M,-2}\colon\mathds{R}(0)\to\mathds{R}(1)$.
@@
The components $\delta_{M,0}$ and $\delta_{M,-1}$ can be more complex, depending on the degree of vanishing of the monodromy operator $N$.
Let $1$ denote the canonical Betti generator of $\mathds{Q}(0)$ and $1^{\vee}$ that of $\mathds{Q}(1)$. Following [Pe06, Theorem 5.19], we define the height jump $\mu\in\mathds{Q}$ as
@@
On the other hand, the limit height is defined by ([BP19, (71)])
@@
Notice that $\mathcal{B}_{\mathcal{H},\lim}$ may not have a biextension structure with respect to the relative weight filtration $M$. The following result relates the height jump and the limit height ([BP19, Theorem 76]).
@@
The amended function $\tilde{h}(t)=h(t)+\mu\log|t|$ extends continuously at $t=0$, with $\tilde{h}(0)=H(N,F_{\infty},W)$.
@@ 4.1. Degeneration of Ceresa cycles
Let $\mathcal{C}\rightarrow\Delta$ be a family as in 1.3.3. We first discuss the case when $C_{0}$ is an irreducible nodal curve, with one node $x_{0}$.
@@
The first statement follows from (1.7), which embeds $C_{0}$ in $X_{0}$, and from (1.4), which implies that the [1][-1] morphism can be lifted to $X_{0}$. Next we show that $Z_{0}$ is homologous to zero in $H^{4}(X_{0};\mathds{Q}(2))$. The cycle class of $Z_{0}$ lands in the weight zero component of $H^{4}(X_{0};\mathds{Q}(2))$ which by equation (1.5) is given by $H^{4}(\widetilde{X}_{0};\mathds{Q}(2))$. Further, since $Z_{0}|_{JC_{0}}=0$ in $\operatorname{CH}^{2}(JC_{0})$, we use the localization exact sequence of Chow groups
@@
In the normalization $\widetilde{X}_{0}$ of $X_{0}$, the cycle $\widetilde{Z}_{0}$ (the normalization of $Z_{0}$) is homologically trivial by the projective bundle formula. Indeed, recall that $\widetilde{X}_{0}=\mathds{P}(\mathcal{O}_{J\widetilde{C}_{0}}\oplus L_{p-q})$ is the projective bundle over $J\widetilde{C}_{0}$. Under the projective bundle isomorphism
@@
the image of the cycle $\widetilde{Z}_{0}$ is given by $(p-q,\widetilde{C}_{0}-[-1]^{\ast}\widetilde{C}_{0})$. Since both the components of this cycle are homologous to zero, so is $\widetilde{Z}_{0}$. Even more is true: Since by Assumption 1.8 the zero cycle pqp-q is $n$-torsion, the class of $\widetilde{Z}_{0}$ is trivial in $\operatorname{CH}^{2}(\widetilde{X}_{0})\otimes\mathds{Q}$.
Even though $X_{0}$ is a singular variety, we show in the next proposition that purity holds for the pair $(X_{0},Z_{0})$. This is probably well known to the experts, but we provide a proof for the sake of completion.
@@
Moreover, $\ker\Big(H^{4}_{|Z_{0}|}(X_{0};\mathds{Q}(2))\to H^{4}(X_{0};\mathds{Q}(2))\Big)\cong\mathds{Q}(0)$, generated by the fundamental class $\{Z_{0}\}$
@@
The fact that $H^{4}_{|Z_{0}|}(X_{0};\mathds{Q}(2))\cong\mathds{Q}(0)\oplus\mathds{Q}(0)$ follows from the standard Verdier duality isomorphism $H^{4}_{|Z_{0}|}(X_{0};\mathds{Q}(2))\cong H^{0}(Z_{0};\mathds{Q}(0))$, using the fact that $|Z_{0}\cap S|=\{(p-b,0),(b-p,\infty)\}$ is finite. Next, for ease of writing, we rename $U=JC_{0}$, which is the open component of $X_{0}$, $S=J\widetilde{C}_{0}$, the singular locus, and $Z_{0,U}\coloneqq Z_{0}\cap U$. We will use the following standard facts: From definition, the cohomology with compact support is given by $H^{k}_{c}(U;\mathds{Q}(2))=H^{k}(X_{0},S;\mathds{Q}(2))$, and $H^{k}_{|Z_{0,U}|,c}(U;\mathds{Q}(2))=H^{k}_{|Z_{0,U}|}(U;\mathds{Q}(2))$. Now, we specialize the long exact sequence of relative cohomology with support in $|Z_{0}|$.
$\cdots\to H^{k-1}_{|Z_{0}\cap S|}(S;\mathds{Q}(2))\to H^{k}_{|Z_{0,U|}}(U;\mathds{Q}(2))\to H^{k}_{|Z_{0}|}(X_{0};\mathds{Q}(2))\to H^{k}_{|Z_{0}\cap S|}(S;\mathds{Q}(2))\to H^{k+1}_{|Z_{0,U}|}(U;\mathds{Q}(2))\to\cdots,$
@@ 4.2. Degeneration of height pairings
In order to state our main result, let us first recall that, in our first degeneration, pulling the cycles $Z_{0},W_{0}$ in $X_{0}$ back to the normalization $\widetilde{X}_{0}$, we get two new cycles $\widetilde{Z}_{0}$ and $\widetilde{W}_{0}$ that intersect properly, and are homologous (in fact rationally equivalent) to zero in $Z^{2}(\widetilde{X}_{0})$. Since $\widetilde{X}_{0}$ is smooth and projective of dimension three, we get another biextension mixed Hodge structure
@@
with associated height $\operatorname{ht}(B_{\widetilde{Z}_{0},\widetilde{W}_{0}})$. In the case of our second degeneration, $X_{0}$ is smooth, and no normalization is needed. Nevertheless, with an abuse of notation, even in this case, we will refer to the normalized biextension in 4.3. The rest of the section is devoted to the following theorem:
@@
Let $\mathcal{C}\to\Delta$ be a family of genus three curves with central fibre $C_{0}$ of the two types discussed above. Let $\mathcal{Z}=\{Z_{t}\}_{t\in\Delta^{\ast}}$ and $\mathcal{W}=\{W_{t}\}_{t\in\Delta^{\ast}}$ be two families of Ceresa cycles obtained by choosing the base points $b_{i,t},i=1,2$ as above. Then we have
$\lim_{t\to 0}h(B_{Z_{t},W_{t}})=\operatorname{ht}(B_{\widetilde{Z}_{0},\widetilde{W}_{0}}).$
@@
By [HR04, Theorem 1.3], the height jump $\mu=0$ for the two cases of our interest. Hence, once we show that $h(B_{\widetilde{Z}_{0},\widetilde{W}_{0}})$ equals the limit height of equation (3.2), Proposition 3.10 would imply the statement of Theorem 4.7. To do this, we first construct the limit mixed Hodge structure associated to the family $\mathcal{V}_{t}\coloneqq H^{3}(JC_{t}\setminus|Z_{t}|,|W_{t}|;\mathds{Q}(2))$ of biextension mixed Hodge structures. Even though conceptually it follows from the functorial properties of limit mixed Hodge structures, we provide the details since it is necessary to obtain our result.
@@
Also, it is evident that the unique generator $1\in\mathds{Q}(0)$ maps to $\{Z_{0}\}$, which in turn maps to $\{\widetilde{Z}_{0}\}$ under $\nu^{\ast}$.
@@
By equation (3.2), $\delta_{M,-2}$ defines the limit height $H(N,F_{\infty},W)$. Next, let us recall that, by Remark 4.2, $\widetilde{Z}_{0},$ and $\widetilde{W}_{0}$, are rationally trivial in $\operatorname{CH}^{2}(\widetilde{X}_{0})\otimes\mathds{Q}$. This implies that the Abel-Jacobi class $\operatorname{AJ}(\widetilde{Z}_{0})\in JH^{3}(\widetilde{X}_{0};\mathds{Q}(2))$ is the splitting extension class
@@
Indeed, the inclusion $H^{1}(C_{0};\mathds{Q})\hookrightarrow H^{1}_{\lim}(C_{t};\mathds{Q})$, and the fact that $H^{3}_{\lim}(JC_{t};\mathds{Q})=\wedge H^{1}_{\lim}(C_{t};\mathds{Q})$, shows that $H^{3}(X_{0};\mathds{Q}(2))\hookrightarrow H^{3}_{\lim}(JC_{t};\mathds{Q}(2))$. Further, using five lemma we obtain
@@
and using (relevant versions of) diagram 1, we conclude that $\iota$ is also injective. Let us denote $\mathcal{Q}=\operatorname{Im}(\iota)\subset\mathcal{V}_{\lim}$. We note that $\operatorname{Gr}_{k}\iota$ induces isomorphism for k=1,2k=-1,-2, while $\operatorname{Gr}_{0}\iota$ is injective. By (4.6), it is also evident that $\operatorname{Gr}_{0}\iota$ maps $\{Z_{0}\}$ to the canonical Betti generator $1\in\mathds{Q}(0)=\operatorname{Gr}^{M}_{0}\mathcal{V}_{\lim}$.
@@
Therefore, if we restrict to the component $\delta^{-1,-1}_{\mathcal{Q}}\colon\mathds{R}(0)\to\mathds{R}(1)$ of $\delta_{\mathcal{Q}}$ and to $\delta_{M,-2}$ of $\delta_{M}$, we read the condition (4.11), as $\delta^{-1,-1}_{\mathcal{Q}}=\delta_{M,-2}$. Hence, we can can restate the limit height formula as
$\delta^{-1,-1}_{\mathcal{Q}}(1)=\frac{1}{2\pi}H(N,F_{\infty},W)1^{\vee},$
where, as before $1\in\mathds{Q}(0)$ and $1^{\vee}\in\mathds{Q}(1)$ are the canonical Betti generators. Now, define

### 15 Apr 2026 v2

@@ 2.2.1. The case of central irreducible nodal curve
We consider the first degeneration described in subsection 1.2.3: Here, the central fibre over $0\in\Delta$ is an irreducible nodal curve $C_{0}$, the node $x_{0}$ is obtained by identifying two points $p,q$ in the normalization $\widetilde{C}_{0}$. Furthermore, the cycle pqp-q is torsion. The natural variation of Hodge structure, whose fibre over $t\in\Delta$ is provided by $H^{1}(C_{t},\mathbb{Q})$, has a limit MHS, as we now recall. Let $H_{1}(C_{t},\mathbb{Z})=\langle\alpha_{1},\alpha_{2},\alpha_{3},\beta_{1},\beta_{2},\beta_{3}\rangle$ be a symplectic basis (similarly $\langle\alpha^{*}_{1},\alpha^{*}_{2},\alpha^{*}_{3},\beta^{*}_{1},\beta^{*}_{2},\beta^{*}_{3}\rangle$ the dual basis in cohomology), and $\alpha_{3}$ the vanishing cycle of the degeneration. Then, since $N^{2}=0$, the monodromy weight filtration has 3 weights and
@@
where the equality in $M_{1}$ follows from Theorem 2.1, since up to a base change, our family admits a semistable degeneration. We want the mixed Hodge structure on $H^{1}_{lim}(C_{t};\mathds{Q})$ to be split, namely we want $H^{1}_{lim}(C_{t};\mathds{Q})=\oplus_{k}\operatorname{Gr}^{M}_{k}H^{1}_{lim}(C_{t};\mathds{Q})$, where
@@
as a split mixed Hodge structure. Now, using the fact that $H^{k}(JC_{t};\mathds{Q})\cong\wedge^{k}H^{1}(C_{t};\mathds{Q})$, we compute $H^{k}_{lim}(JC_{t}:\mathbb{Q})$, by taking the exterior power $\wedge^{k}H^{1}_{lim}(JC_{t};\mathbb{Q})\cong\wedge^{k}H^{1}_{lim}(C_{t};\mathbb{Q})$. We keep calling the nilpotent logarithm of monodromy $\wedge^{k}N$ of the variation $H^{k}(JC_{t};\mathds{Q})$ by $N$, satisfying $N^{2}=0$.
It follows that the limit MHS $H^{k}_{lim}(JC_{t};\mathbb{Q})$ has 3 weights with respect to relative weight filtration $M$, with graded pieces:
$\operatorname{Gr}^{M}_{k-1}H^{k}_{lim}(JC_{t};\mathbb{Q})=H^{k-1}(J\widetilde{C}_{0},\mathds{Q}),$
$\operatorname{Gr}^{M}_{k}H^{k}_{lim}(JC_{t};\mathbb{Q})=H^{k-2}(J\widetilde{C}_{0};\mathds{Q}(-1))\oplus H^{k}(J\widetilde{C}_{0};\mathds{Q}))\cong H^{k}(\widetilde{X}_{0};\mathds{Q}),$
$\operatorname{Gr}^{M}_{k+1}H^{k}_{lim}(JC_{t};\mathbb{Q})=H^{k-1}(J\widetilde{C}_{0};\mathds{Q}(-1)).$
The isomorphism in the middle follows from the projective bundle formula, noting that $\widetilde{X}_{0}$ is $\mathbb{P}^{1}$-bundle over $J\widetilde{C}_{0}$.
@@ 2.2.2. The case of central reducible curve
We now consider the second degeneration in 1.2.3: Here the central fibre over $0\in\Delta$ is a reducible curve $C_{0}=C_{1}\cup C_{2}$ obtained as the union of two irreducible smooth components $C_{1}$ and $C_{2}$ of genera $1$ and $2$ respectively, with a separating node ${y_{0}}=C_{1}\cap C_{2}$. As before, we first consider the VHS provided by $H^{1}(C_{t};\mathds{Q})$, for $t\in\Delta^{*}$. In this case, $N=0$ and the limit MHS trivially is as follows:
@@
Remark 2.3.
Since $N=0$, this is the case for no degeneration, and we technically do not need assumption 2.2. However, we work with it to bring both the degenerations under a common set up.
@@ 3.1. Geometric biextension
In case of a biextension of geometric origin as above, the corresponding height $\operatorname{ht}(B_{Z,W})$ equals the archimedean height pairing $\langle Z,W\rangle$ between the two cycles (this is a consequence of [Ha90, Proposition 3.2.13]. See also [Pe06, §5] for further insights.). In fact, using the duality of remark 3.5 one can interpret $\operatorname{ht}(B_{Z,W})$ as $\int_{Z}\mathfrak{g}_{W}$, where ${\mathfrak{g}}_{W}$ is a Green form of logarithmic type for the cycle $W$ that can be shown to exist. For example, in $\mathds{P}^{1}$ any choice of four distinct points $\{p,q,r,s\}$ gives rise to homologically (actually rationally) trivial cycles Z=pqZ=p-q and W=rsW=r-s. In this case, one choice for ${\mathfrak{g}}_{W}$ is the function $\log\Big|\frac{t-r}{t-s}\Big|^{2}$ that has logarithmic singularities on $|W|=\{r,s\}$. The height is given by the logarithm of the cross-ratio of the four points.
@@ 3.2. Abstract degeneration of biextension variations
This function does not always extend continuously to $\Delta$. There is a $\mu\in\mathds{Q}$, such that the amended function
@@
with respect to the original weight filtration $W$. Hence the components $N_{-2}$ are $\delta_{M,-2}$ are easy to extract: $N_{-2}$ is the component of $N$ that goes from $\mathds{Q}(0)$ to $\mathds{Q}(1)$, while $\delta_{M,-2}$ is the component of $\delta_{M}$ that goes from $\mathds{R}(0)$ to $\mathds{R}(1)$.
@@
Notice that, $\mathcal{B}_{\mathcal{H},\lim}$ may not have a biextension structure with respect to the relative weight filtration $M$. Also, the components $\delta_{M,0}$ and $\delta_{M,-1}$ can be more complex, depending on the degree of vanishing of the monodromy operator $N$.
Let $1$ denote the canonical Betti generator of $\mathds{Q}(0)$ and $1^{\vee}$ that of $\mathds{Q}(1)$. Following [Pe06, Theorem 5.19], we define $\mu$ by
@@
Since $N$ does not depend on the choice of a coordinate for $\Delta$, neither does $\mu$. On the other hand, the limit height is defined by ([BP19, (71)])
@@
In this case, since the limit mixed Hodge structure depends on a choice of coordinate for $\Delta$, so does the height $H(N,F_{\infty},W)$. The following result justifies the name ‘limit height’ ([BP19, Theorem 76]).
@@
The asymptotic height $\tilde{h}(t)=h(t)+\mu\log|t|$ extends continuously to $t=0$, with $\tilde{h}(0)=H(N,F_{\infty},W)$.
@@ 4.1. Degeneration of Ceresa cycles
Let $\mathcal{C}\rightarrow\Delta$ be a family as in 1.2.3. We first discuss the case when $C_{0}$ is an irreducible nodal curve, with one node $x_{0}$.
@@
The first statement follows from (1.7), which embeds $C_{0}$ in $X_{0}$, and from (1.4), which implies that the [1][-1] morphism can be lifted to $X_{0}$. Next, using standard Verdier duality isomorphism $H^{4}_{|Z_{0}|}(X_{0};\mathds{Q}(2))\cong H^{0}(Z_{0};\mathds{Q}(0))$ and the fact that $|Z_{0}\cap S|=\{(p-b,0),(b-p,\infty)\}$ is finite, we get that $H^{4}_{|Z_{0}|}(X_{0};\mathds{Q}(2))\cong\mathds{Q}(0)\oplus\mathds{Q}(0)$. The morphism $H^{4}_{|Z_{0}|}(X_{0};\mathds{Q}(2))\to H^{4}(X_{0};\mathds{Q}(2))$ defines the cycle class map. We show that $Z_{0}$ is homologous to zero in $H^{4}(X_{0};\mathds{Q}(2))$. The cycle class of $Z_{0}$ lands in the weight zero component of $H^{4}(X_{0};\mathds{Q}(2))$ which by equation (1.6) is given by $H^{4}(\widetilde{X}_{0};\mathds{Q}(2))$. Further, since $Z_{0}|_{JC_{0}}=0$ in $\operatorname{CH}^{2}(JC_{0})$, we use the localization exact sequence of Chow groups
@@
In the normalization $\widetilde{X}_{0}$ of $X_{0}$, the class of the cycle $\widetilde{Z}_{0}$ (the normalization of $Z_{0}$) has an explicit description using the projective bundle formula. Indeed, recall that $\widetilde{X}_{0}=\mathds{P}(\mathcal{O}_{J\widetilde{C}_{0}}\oplus L_{p-q})$ is the projective bundle over $J\widetilde{C}_{0}$. Under the projective bundle isomorphism
@@
the image of the cycle $\widetilde{Z}_{0}$ is given by $(p-q,\widetilde{C}_{0}-[-1]^{\ast}\widetilde{C}_{0})$, where we view pqp-q as a zero cycle in $J\widetilde{C}_{0}$. Since by assumption 1.8 the zero cycle pqp-q is torsion, the class of $\widetilde{Z}_{0}$ is trivial in $\operatorname{CH}^{2}(\widetilde{X}_{0})\otimes\mathds{Q}$. Hence $\widetilde{Z}_{0}$ is rationally equivalent to zero in $Z^{2}(\widetilde{X}_{0})\otimes\mathds{Q}$.
Even though $X_{0}$ is a singular variety, we show in the next proposition that purity holds for the pair $(X_{0},Z_{0})$. This is probably well known to experts, but we provide a proof for completeness.
@@
Moreover, $\ker\Big(H^{4}_{|Z_{0}|}(X_{0};\mathds{Q}(2))\to H^{4}(X_{0};\mathds{Q}(2))\Big)\cong\mathds{Q}(0)$, generated by the fundamental class $\{Z_{0}\}$.
@@
The equality for $k=4$ is already shown in Proposition 4.1. For ease of writing, we rename $U=JC_{0}$, which is the open component of $X_{0}$, $S=J\widetilde{C}_{0}$, the singular locus, and $Z_{0,U}\coloneqq Z_{0}\cap U$. We will use the following standard facts: From definition, the cohomology with compact support is given by $H^{k}_{c}(U;\mathds{Q}(2))=H^{k}(X_{0},S;\mathds{Q}(2))$, and $H^{k}_{|Z_{0,U}|,c}(U;\mathds{Q}(2))=H^{k}_{|Z_{0,U}|}(U;\mathds{Q}(2))$. Now, we specialize the long exact sequence of relative cohomology with support in $|Z_{0}|$
$\cdots\to H^{k-1}_{|Z_{0}\cap S|}(S;\mathds{Q}(2))\to H^{k}_{|Z_{0,U|}}(U;\mathds{Q}(2))\to H^{k}_{|Z_{0}|}(X_{0};\mathds{Q}(2))\\ \to H^{k}_{|Z_{0}\cap S|}(S;\mathds{Q}(2))\to H^{k+1}_{|Z_{0,U}|}(U;\mathds{Q}(2))\to\cdots,$
@@ 4.2. Degeneration of height pairings
To state our main result, let us first recall that, in our first degeneration, pulling the cycles $Z_{0},W_{0}$ in $X_{0}$ back to the normalization $\widetilde{X}_{0}$, we get two new cycles $\widetilde{Z}_{0}$ and $\widetilde{W}_{0}$ that still intersect properly, and are rationally equivalent to zero in $Z^{2}(\widetilde{X}_{0})\otimes\mathds{Q}$. Since $\widetilde{X}_{0}$ is smooth and projective of dimension three, we get another biextension mixed Hodge structure
@@
with associated height $\operatorname{ht}(B_{\widetilde{Z}_{0},\widetilde{W}_{0}})$. In the case of our second degeneration, $X_{0}$ is smooth, and no normalization is necessary. The rest of the section is devoted to the following theorem:
@@
Let $\mathcal{C}\to\Delta$ be a family of genus three curves with central fibre $C_{0}$ of the two types discussed above. Let $\mathcal{Z}=\{Z_{t}\}_{t\in\Delta^{\ast}}$ and $\mathcal{W}=\{W_{t}\}_{t\in\Delta^{\ast}}$ be two families of Ceresa cycles obtained by choosing the base points $b_{i,t},i=1,2$ as above. Then, for a fixed choice of coordinate (assumption 2.2), the limit height $H(N,F_{\infty},W)$ associated with the biextension variation $B_{Z_{t},W_{t}}$ is given by the Deligne splitting of $B_{\widetilde{Z}_{0},\widetilde{W}_{0}}$. Consequently, we have $H(N,F_{\infty},W)=\operatorname{ht}(B_{\widetilde{Z}_{0},\widetilde{W}_{0}})$.
@@
The key point is to construct (under the assumptions 1.8 and 2.2) the limit mixed Hodge structure associated with the variation $\mathcal{V}_{t}\coloneqq H^{3}(JC_{t}\setminus|Z_{t}|,|W_{t}|;\mathds{Q}(2))$ of biextension mixed Hodge structures. Even though conceptually it follows from the functorial properties of limit mixed Hodge structures, we provide the details for clarity.
@@
Also, it is evident that the unique generator $1_{t}\in\mathds{Q}(0)$ maps to $\{Z_{0}\}$, which in turn maps to $\{\widetilde{Z}_{0}\}$ under $\nu^{\ast}$. The subscript $t$ reflects the choice of coordinate made in assumption 2.2.
@@
By equation (3.2), $\delta_{M,-2}$ defines the limit height $H(N,F_{\infty},W)$. Next, let us recall that, by Remark 4.2, $\widetilde{Z}_{0},$ and $\widetilde{W}_{0}$, are rationally trivial in $Z^{2}(\widetilde{X}_{0})\otimes\mathds{Q}$. This implies that the Abel-Jacobi class $\operatorname{AJ}(\widetilde{Z}_{0})\in JH^{3}(\widetilde{X}_{0};\mathds{Q}(2))$ is the splitting extension class
@@
Indeed, the inclusion $H^{1}(C_{0};\mathds{Q})\hookrightarrow H^{1}_{\lim}(C_{t};\mathds{Q})$, and the fact that $H^{3}_{\lim}(JC_{t};\mathds{Q})=\wedge H^{1}_{\lim}(C_{t};\mathds{Q})$, shows that $H^{3}(X_{0};\mathds{Q}(2))\hookrightarrow H^{3}_{\lim}(JC_{t};\mathds{Q}(2))$. Further, using the five lemma, we obtain
@@
and using (relevant versions of) diagram 1, we conclude that $\iota$ is also injective. Let us denote $\mathcal{Q}=\operatorname{Im}(\iota)\subset\mathcal{V}_{\lim}$. We note that $\operatorname{Gr}_{k}\iota$ induces isomorphism for k=1,2k=-1,-2, while $\operatorname{Gr}_{0}\iota$ is injective. By (4.6), it is also evident that $\operatorname{Gr}_{0}\iota$ maps $\{Z_{0}\}$ to the canonical Betti generator $1_{t}\in\mathds{Q}(0)=\operatorname{Gr}^{M}_{0}\mathcal{V}_{\lim}$.
@@
Therefore, if we restrict to the component $\delta^{-1,-1}_{\mathcal{Q}}\colon\mathds{R}(0)\to\mathds{R}(1)$ of $\delta_{\mathcal{Q}}$ and to $\delta_{M,-2}$ of $\delta_{M}$, we read the condition (4.11), as $\delta^{-1,-1}_{\mathcal{Q}}=\delta_{M,-2}$. Hence, we can restate the limit height formula as
$\delta^{-1,-1}_{\mathcal{Q}}(1_{t})=\frac{1}{2\pi}H(N,F_{\infty},W)1^{\vee}_{t},$
where, as before $1_{t}\in\mathds{Q}(0)$ and $1^{\vee}_{t}\in\mathds{Q}(1)$ are the canonical Betti generators. Now, define